# Neutrino Decoherence via Modified Dispersion

### v1

@@ 5 Energy Dependence of Decoherence
where $E_{0}$ is a constant energy scale which we set to $E_{0}=1$ GeV, and $n$ is a power-law index to be tested experimentally. The parameter $n$ can take both positive and negative values, including zero. We discuss a few implications depending on this as follows.
@@
    Lightcone Fluctuating Model : ($n=-2$)
@@
    Lorentz Invariant Model : ($n=-1$)
@@ 6.1 Observables and Relevant Parameters
    Decoherence Strength or Relaxation: Three independent strength may arise from this model, i.e., $\Gamma_{21}(E)$, $\Gamma_{31}(E)$ and $\Gamma_{32}(E)$. These can provide us with new constraints on the upper limits and also help measure the $\delta_{CP}$ phase.
@@ 6.2 Decoherence Strength Parameters
It is evident from the previous expression, different values of $a$ correspond to different energy dependence ($n=2a-2$) which has been briefly explained in Sec. 5.
@@ 6.3 Coherence Length
In the case of solar neutrinos with energies $\mathcal{O}(10^{-4})$ GeV, coherence lengths clearly follow a trend akin to that of the cosmic neutrino background (CNB) neutrinos. Decoherence signatures become pronounced as the parameter $a$ changes from 00 to $1$, even at these lower energy levels. In parallel, for atmospheric neutrinos with energies $\mathcal{O}(10^{3})$ GeV, the impact of decoherence manifests over distances well below the Earth’s diameter, remaining observable till the propagation length is comparable to the radius of the observable universe for all possible values of $a$. Furthermore, for ultra high energy (UHE) neutrinos of energy $\mathcal{O}(10^{6})$ GeV, it can be seen that only near to the Earth-Sun distance, the neutrinos lose coherence. As $a\rightarrow 1$, the effects become more prominent, making it verifiable in observatories on Earth.
@@ 6.4 Detection Prospects
    Earth Diameter: For a propagation length of about $10^{4}$ km, achieving the detectable sensitivity bound for the decoherence strength parameter requires an energy threshold of around $10^{3}$ GeV. This energy level aligns well with the detection capabilities of experiments such as IceCube and its future generation counterpart, IceCube Gen-2.
@@
    Earth-Sun Distance: To meet the detectable sensitivity bound for the decoherence strength parameter across a propagation length of roughly $10^{8}$ km, an energy threshold of about $10^{5}$ GeV is essential. Experiments such as IceCube Gen-2 and ANITA possess the necessary detection capabilities to explore these energies.
@@
    Milky way Diameter: In order to attain the detectable sensitivity bound for the decoherence strength parameter over a propagation length of roughly $10^{18}$ km, an energy threshold of approximately $10^{9}$ GeV is imperative. Experiments such as GRAND and RNO-G are equipped with the sophisticated detection capabilities necessary to investigate these energy levels.
@@
    Radius of Observable Universe: Similarly, for a baseline of $10^{24}$ km, to meet the detectable sensitivity bound for the decoherence strength parameter requires an energy threshold of around $10^{12}$ GeV. Experiments such as the proposed POEMMA, ARIANNA are projected to be able to meet such energy levels.
@@ 6.5 Oscillation Probabilities
In the top left panel of Fig. 3, the muon neutrino survival probability across different propagation lengths is shown. Here, we show the neutrino oscillation after different propagation lengths in a fixed interval. In each case we illustrate the neutrino oscillation probabilities between the propagation lengths of $L^{\prime}$ and $L$ such that the interval $(L-L^{\prime})\in[10^{14},6\times 10^{14}]\mathrm{~{}km}$. It is to be noted here that depicting the oscillation probabilities in the same interval after different propagation length is one of the most efficient way of presenting how the decohrence affects the neutrino oscillation. Notably, as we increase the neutrino propagation length $L^{\prime}$ given to the neutrino, the damping signatures on the probabilities become strikingly evident. Similarly, on the top right of the panel, the $\nu_{\mu}\rightarrow\nu_{\tau}$ transition probability also highlights the extensive impact of decoherence. The effects for the tau-neutrino family are effectively summarized bottom-left and bottom-right panels of Fig. 3 respectively. Importantly, as the distance travelled by the neutrino increases, we observe the oscillatory (“wiggly”) patterns diminishing, highlighting the substantial influence of decoherence in this context.
@@
[…]
@@
On the top of Fig. 5, the electron neutrino survival probability is plotted against the propagation length, varying the parameter $a$ from 00 to $1$. In particular, we have chosen $a=0,0.1,1$. In the lower propagation length region and for $a\rightarrow 1$, the decoherence effect is almost negligible. However, as $a\rightarrow 0$, the influence of decoherence becomes more pronounced as the oscillatory behaviour dies out, even at these lower baselines, till the equilibration happens. This highlights the fascinating relationship between decoherence and oscillation probabilities.
On the bottom-left and bottom-right of Fig. 5, the muon neutrino survival probabilities further illustrate this trend. For smaller values of $a$, we can observe a notable impact due to decoherence, which adds depth to our analysis. Moreover, within the propagation length range of $10^{9}$ km to $10^{11}$ km, a smooth transition emerges as a direct result of decoherence effects, suggesting a complex yet intriguing dynamic at play. Interestingly, beyond $a=1$ and below $a=0$, the observed behaviour continues to show an insignificant impact of decoherence on all probabilities, providing an avenue for constraining the parameter $a$ and understanding its role. A similar trend can be seen on the bottom-right of Fig. 5.
@@ 6.6 Flavour Composition at Neutrino Telescopes
As seen from the left of Fig. 6, the flavour compositions of neutrinos change as they travel from their source to Earth. The expected ratios for three different pre-oscillation initial abundances of (1 : 2 : 0), (1 : 1 : 0), (1 : 0 : 0), and (0 : 1 : 0) are represented by green, red, blue, and magenta markers, respectively. At Earth-based neutrino telescopes, the measured post-oscillation abundances, encircled by a yellow marker, are (0.38 : 0.31 : 0.31), (0.34 : 0.34 : 0.32), (0.25 : 0.43 : 0.32), and (0.43 : 0.24 : 0.23). However, when incorporating decoherence arising due to MDR into the analysis alongside post-oscillation data, the measurements shift to (0.35 : 0.32 : 0.32), (0.41 : 0.31 : 0.28), (0.55 : 0.26 : 0.19), and (0.26 : 0.36 : 0.38), respectively. This indicates a significant and verifiable influence of decoherence on the results. All these measured values correspond to normal hierarchy (NH). While on the right of Fig. 6, we conducted similar measurements with inverted hierarchy (IH). The post-oscillation observations reveal the flavour compositions as follows: (0.34 : 0.33 : 0.33), (0.42 : 0.29 : 0.29), (0.65 : 0.17 : 0.15), and (0.17 : 0.41 : 0.42) for the four outlined sources detailed in Sec . 6.6. Notably, when decoherence is incorporated into the analysis, the observed flavour compositions shift to (0.32 : 0.35 : 0.32), (0.38 : 0.30 : 0.32 (0.55 : 0.23 : 0.22), and (0.23 : 0.39 : 0.38). These results underscore significant distinctions in measurements across the two different hierarchies. This signature of decoherence on neutrino oscillation properties decisively opens new avenues to address the still-open question of neutrino preferred hierarchy.
@@ 7 Extension to ($3+1$) flavour scenario
[…]
@@
$\pi$ decay 1 2 0 0 0.34 0.29 0.28 0.09 0.33 0.30 0.26 0.11
$c$ decay 1 1 0 0 0.51 0.21 0.19 0.09 0.46 0.25 0.21 0.08
$n$ decay 1 0 0 0 0.96 0.01 0.01 0.02 0.85 0.08 0.05 0.02
$\mu$ damped 0 1 0 0 0.01 0.42 0.44 0.13 0.10 0.40 0.36 0.14
Table 3: Here, we show initial ($\Phi^{i}_{\alpha}$) and final ($\Phi^{f}_{\alpha}$) neutrino flavour compositions for various production mechanisms, with and without decoherence effects in $3+1$ formalism varying $a$ from 00 to $1$ and keeping the neutrino energy $1$ PeV and baseline of $10^{12}$ km.
The flavour compositions of the neutrinos $\nu_{e}$, $\nu_{\mu}$, $\nu_{\tau}$, and $\nu_{s}$ are clearly illustrated in a normalized form in Fig. 8 and Table  3. The post-oscillation observations, which exclude decoherence effects and assume $\nu_{s}=0$ at the source, deliver the compositions: (0.34 : 0.29 : 0.28 : 0.09), (0.51 : 0.21 : 0.19 : 0.09), (0.96 : 0.01 : 0.01 : 0.02), and (0.01 : 0.42 : 0.44 : 0.13) for the four initial compositions outlined in Sec. 6.6, respectively. However, when we incorporate the effects of decoherence, the data showcases a significant shift in these values. The revised compositions are (0.33 : 0.30 : 0.26 : 0.11), (0.46 : 0.25 : 0.21 : 0.08), (0.85 : 0.08 : 0.05 : 0.02), and (0.10 : 0.40 : 0.36 : 0.14), showcasing finite differences in both cases. This evidence decisively indicates that the inclusion of heavy sterile neutrinos has a profound impact on flavour composition measurements on Earth, and advancing decoherence studies will undoubtedly yield critical insights into these scenarios.
@@ 8 Summary and Conclusion
After briefly discussing the existing treatments of neutrino oscillations, we show the mathematical formalism of describing neutrinos as an open quantum system, where we discuss the possibility of neutrino decoherence in the three and $3+1$ flavour scenarios. We explain the modified dispersion relation and discuss the relevant parameters that contain the vital information regarding the various quantum gravity theories. In this regard, it is to be noted that in order to relate the modified dispersion relation with existing theories, we use a variable $a$ which can take values between 00 and $1$, where $a=1$ denotes that the modification to the dispersion relation is Planck energy suppressed, whereas $a=0$ signifies that the modification is entirely dependent on the energy of the system.
@@
We consider the decoherence parameter (and coherence length), oscillation probability, and the flux composition as the main observables. We have shown that for $a=1$ the decoherence parameter and coherence length are independent of neutrino energy, whereas for $a=0$ the decoherence parameter (coherence length) decreases (increases) with energy. Next we show the evolution of the oscillation probability in this decoherence setup in two different ways. In the first case, we show how the oscillation probabilities change with propagation length and we see that for $a=2/3$ and neutrino energy of $10^{8}$ GeV, the oscillation probabilities show a significant change after propagating large distances. In order to understand this behaviour further, we also show the oscillation probabilities averaged over propagation lengths for both normal and inverted hierarchy of neutrino masses. We see that the effect of decoherence manifests differently in normal and inverted hierarchy cases. We then show the evolution of the oscillation probabilities for a $100$ PeV neutrino with in a propagation length ranging from $10^{9}$-$10^{15}$ km and constraining $a$ between 00 and $1$. As expected, the decoherence sets in at very low distances for $a=0$, whereas only for $a=1$ the decoherence is negligible for low energies, but for high energies the decoherence becomes dominant, where the survival (transition) probability sees a drop (rise). Finally, we show the flux composition reaching the Earth and the effect of decoherence in that aspect. We see that the effect of decoherence in this case depends on the initial flux composition, the $a$ value, and the mass hierarchy, i.e., for inverted hierarchy, the deviation of the final flux composition due to decoherence is considerably different than the normal hierarchy.
@@
Apart from the three-flavour case (as discussed above), we also show the effect of decoherence on the oscillation probabilities for a $3+1$ scenario. We show that, however, overall the features of the evolution of survival and transition properties look similar to the three-flavour case, the dip (rise) in the survival (transition) probability has an anomalous drop in the $3+1$ flavour case. This drop can be directly linked to the existence of the extra flavour. On the other hand, due to the presence of the extra flavour, the final flux composition is different in the $3+1$ flavour case than that of the three flavour case.

### v2

@@ 5 Energy Dependence of Decoherence
where $E_{0}$ is a constant energy scale which we set to $E_{0}=1$ GeV, and $n$ is a power-law index to be tested experimentally. While $\Gamma_{ij}$ appears time-independent in Eq. (5.1), it is important to note that an explicit time dependence may emerge if the environment possesses stochastic or dynamical properties, as determined by the specific model under consideration. The parameter $n$ can take both positive and negative values, including zero. We discuss a few implications depending on this as follows.
@@
    Lightcone Fluctuating Model: (n=2n=-2)
@@
    Lorentz Invariant Model: (n=1n=-1)
@@ 6.1 Observables and Relevant Parameters
    Decoherence Strength or Relaxation: Three independent strength may arise from this model, i.e., $\Gamma_{21}(E)$, $\Gamma_{31}(E)$ and $\Gamma_{32}(E)$. As a result, different modifications of oscillation patterns will occur after different length of propagation which can in principle be probed by forthcoming experiments.
@@ 6.2 Decoherence Strength Parameters
It is evident from the previous expression, different values of $a$ correspond to different energy dependence (n=2a2n=2a-2) which has been briefly explained in Sec. 5. Also, it is to be noted that, once the neutrino source is fixed, the dissipator effectively becomes time (or length)-independent.
@@ 6.3 Coherence Length
In the case of solar neutrinos with energies $\mathcal{O}(10^{-4})$ GeV, coherence lengths clearly follow a trend akin to that of the cosmic neutrino background (CNB) neutrinos. Decoherence signatures become pronounced as the parameter $a$ changes from 0 to $1$, even at these lower energy levels. In parallel, for atmospheric neutrinos with energies $\mathcal{O}(10^{3})$ GeV, the impact of decoherence manifests over distances well below the Earth’s diameter, remaining observable till the propagation length is comparable to the radius of the observable universe for all possible values of $a$. Furthermore, for ultra high energy (UHE) neutrinos of energy $\mathcal{O}(10^{6})$ GeV, it can be seen that only near to the Earth-Sun distance, the neutrinos lose coherence. As $a\rightarrow 1$, the effects become more prominent, making it verifiable in observatories on Earth.
Here we note that the MSW effects [89, 90, 91], which is relevant for neutrinos moving in matter with varying density, may not be significant here since irrespective of the source, e.g., solar or supernova, the neutrinos are going to cover enormous distances to accumulate quantum gravitational effects proportionately. Therefore, for most part of the trajectory it will resemble vacuum oscillations and thus we are neglecting the matter effects in our study.
@@ 6.4 Detection Prospects
    Earth Diameter: For a propagation length of about $10^{4}$ km, achieving the detectable sensitivity bound for the decoherence strength parameter requires an energy threshold of around $10^{3}$ GeV. This energy level aligns well with the detection capabilities of experiments such as IceCube [92] and its future generation counterpart, IceCube Gen-2 [28].
@@
    Earth-Sun Distance: To meet the detectable sensitivity bound for the decoherence strength parameter across a propagation length of roughly $10^{8}$ km, an energy threshold of about $10^{5}$ GeV is essential. Experiments such as KM3NeT [93], IceCube Gen-2 [28] and ANITA [94] possess the necessary detection capabilities to explore these energies.
@@
    Milky way Diameter: In order to attain the detectable sensitivity bound for the decoherence strength parameter over a propagation length of roughly $10^{18}$ km, an energy threshold of approximately $10^{9}$ GeV is imperative. Experiments such as GRAND [35] and RNO-G [19] are equipped with the sophisticated detection capabilities necessary to investigate these energy levels.
@@
    Radius of Observable Universe: Similarly, for a baseline of $10^{24}$ km, to meet the detectable sensitivity bound for the decoherence strength parameter requires an energy threshold of around $10^{12}$ GeV. Experiments such as the proposed POEMMA [95], ARIANNA [96] are projected to be able to meet such energy levels.
At this point, we note that high-energy neutrinos that align with the above mentioned coherence lengths can indeed be generated in powerful cosmic accelerators and other sources. These include Active Galactic Nuclei (AGN), interactions between cosmic rays and cosmic microwave background (CMB) photons, particularly around the Greisen-Zatsepin-Kuzmin (GZK) cutoff, quenched superradiance of black holes, and the later stages of Hawking evaporation of Primordial Black Holes (PBHs), especially those found in the tail of the power law mass distribution [97, 98].
@@ 6.5 Oscillation Probabilities
In Fig. 3, we show how the decoherence modifies the neutrino oscillation probabilities after different propagation lengths. It is to be noted carefully that we show the oscillation probabilities between the propagation lengths $L^{\prime}$ and $L$, i.e., we start ‘looking’ at the probabilities after the neutrinos have propagated a length $L^{\prime}$ and ‘look’ at it till the neutrinos propagate a distance $L$. This allows us to observe the neutrino oscillation probability in an interval of $L-L^{\prime}$. Furthermore, we consider $L-L^{\prime}\in[10^{14},6\times 10^{14}]\mathrm{~km}$ in order to observe only a few cycles of the neutrino oscillations. We do this to show how the neutrino oscillation probabilities change in a fixed interval after various propagation lengths. It is to be noted here that depicting the oscillation probabilities in the same interval after different propagation length is one of the most efficient way of presenting how the decohrence affects the neutrino oscillation.
In the top left panel of Fig. 3, the muon neutrino survival probability across different propagation lengths is shown. Notably, as we increase the neutrino propagation length $L^{\prime}$ given to the neutrino, the damping signatures on the probabilities become strikingly evident. Similarly, on the top right of the panel, the $\nu_{\mu}\rightarrow\nu_{\tau}$ transition probability also highlights the extensive impact of decoherence. The effects for the tau-neutrino family are effectively summarized bottom-left and bottom-right panels of Fig. 3 respectively. Importantly, as the distance travelled by the neutrino increases, we observe the oscillatory (“wiggly”) patterns diminishing, highlighting the substantial influence of decoherence in this context.
@@
In Fig. 4, we show the average oscillation probabilities over the propagation length. To be precise, we show the average oscillation probability integrated over the propagation length $L$ and present them for a few values of $L$. The idea behind this is to take away the oscillatory nature of the probability and focus on the amplitude of the same. Although this method of presentation still contains some ‘wigglyness’ in the average probability ($P^{\mathrm{avg}}_{\alpha\beta}$), it also enables us to focus on how the decoherence sets in and diminishes the amplitude of transition/survival probabilities.
As illustrated in the top left panel, the effect of decoherence is notably different between the two mass hierarchies NH and IH. After travelling a distance of approximately $10^{10}$ km, the probability points begin to differ for different hierarchies. Similar characteristics for $\nu_{\mu}\rightarrow\nu_{\tau}$ transition probability is advocated by the plot on the top right panel. This significant divergence provides a crucial opportunity to tackle the long-standing neutrino hierarchy problem. The difference of probabilities unequivocally underlines the necessity of identifying and investigating sources at these specific distances. By strategically concentrating our efforts on these distances, the open hierarchy problem can be decisively addressed. Plots on the bottom left and bottom right of Fig. 4 reinforce these important characteristics.
@@
On the top of Fig. 5, the electron neutrino survival probability is plotted against the propagation length, varying the parameter $a$ from 0 to $1$. In particular, we have chosen $a=0,0.1,1$. In the lower propagation length region and for $a\rightarrow 1$, the decoherence effect is almost negligible. However, as $a\rightarrow 0$, the influence of decoherence becomes more pronounced as the oscillatory behaviour dies out, even at these lower baselines, till the equilibration happens. This highlights the fascinating relationship between decoherence and oscillation probabilities.
On the bottom-left and bottom-right of Fig. 5, the muon and tau neutrino survival probabilities further illustrate this trend. For smaller values of $a$, we can observe a notable impact due to decoherence, which adds depth to our analysis. Moreover, within the propagation length range of $10^{9}$ km to $10^{11}$ km, a smooth transition emerges as a direct result of decoherence effects, suggesting a complex yet intriguing dynamic at play. Interestingly, beyond $a=1$ and below $a=0$, the observed behaviour continues to show an insignificant impact of decoherence on all probabilities, providing an avenue for constraining the parameter $a$ and understanding its role. A similar trend can be seen on the bottom-right of Fig. 5.
@@ 6.6 Flavour Composition at Neutrino Telescopes
As seen from the left of Fig. 6 and Tab. 2, the flavour compositions of neutrinos change as they travel from their source to Earth. The expected ratios for three different pre-oscillation initial abundances of (1 : 2 : 0), (1 : 1 : 0), (1 : 0 : 0), and (0 : 1 : 0) are represented by green, red, blue, and magenta markers, respectively. At Earth-based neutrino telescopes, the measured post-oscillation abundances, encircled by a yellow marker, are (0.38 : 0.31 : 0.31), (0.34 : 0.34 : 0.32), (0.25 : 0.43 : 0.32), and (0.43 : 0.24 : 0.23). However, when incorporating decoherence arising due to MDR into the analysis alongside post-oscillation data, the measurements shift to (0.35 : 0.32 : 0.32), (0.41 : 0.31 : 0.28), (0.55 : 0.26 : 0.19), and (0.26 : 0.36 : 0.38), respectively. This indicates a significant and verifiable influence of decoherence on the results. All these measured values correspond to normal hierarchy (NH). While on the right of Fig. 6, we conducted similar measurements and measured values are given in Tab. 3 with inverted hierarchy (IH). The post-oscillation observations reveal the flavour compositions as follows: (0.34 : 0.33 : 0.33), (0.42 : 0.29 : 0.29), (0.65 : 0.17 : 0.15), and (0.17 : 0.41 : 0.42) for the four outlined sources detailed in Sec . 6.6. Notably, when decoherence is incorporated into the analysis, the observed flavour compositions shift to (0.32 : 0.35 : 0.32), (0.38 : 0.30 : 0.32 (0.55 : 0.23 : 0.22), and (0.23 : 0.39 : 0.38). These results underscore significant distinctions in measurements across the two different hierarchies. This signature of decoherence on neutrino oscillation properties decisively opens new avenues to address the still-open question of neutrino preferred hierarchy. At this point, it is important to acknowledge that, in all the flux composition measurements conducted here, we have chosen to set aside the systematic uncertainties involved.
@@ 7 Extension to ($3+1$) flavour scenario
where $R(\theta_{ij},\delta_{ij})$ is a rotation in the $ij$-th plane with an associated phase $\delta_{ij}$. In this work, we assume all the CP phases [^4] Similar study with finite phases have been carried out in  [54] in a three flavour scenario discriminating Dirac and Majorana phases. $\delta_{ij}=0$ and take the mixing parameter values as mentioned in the Tab. 4 given below.
Parameter Best-fit values
$\sin^{2}\theta_{12}$ $0.317$
$\sin^{2}\theta_{13}$ $0.022$
$\sin^{2}\theta_{14}$ $0.009$
$\sin^{2}\theta_{23}$ $0.566$
$\sin^{2}\theta_{24}$ $0.007$
$\sin^{2}\theta_{34}$ $0.116$
$\Delta m^{2}_{41}$ $\mathrm{eV}^{2}$] $1.69$
Table 4: $3+1$ neutrino oscillation parameters used for numerical analyses in our study taken from Refs. [112, 113].
As mentioned earlier in Sec. 3, here, in this case we can span the needed observables with $4\times 4$ generalized Gell-Mann matrices and can repeat a similar procedure as outlined in Sec. 3 to obtain the MDR-induced master equation in $3+1$ scenario and consequently the relevant observables.
@@
$\pi$ decay 1 2 0 0 0.33 0.34 0.25 0.08 0.34 0.34 0.24 0.08
$c$ decay 1 1 0 0 0.43 0.27 0.24 0.05 0.40 0.30 0.22 0.08
$n$ decay 1 0 0 0 0.65 0.09 0.22 0.06 0.58 0.19 0.19 0.04
$\mu$ damped 0 1 0 0 0.13 0.51 0.27 0.09 0.20 0.42 0.27 0.09
Table 5: Here, we show initial ($\Phi^{i}_{\alpha}$) and final ($\Phi^{f}_{\alpha}$) neutrino flavour compositions for various production mechanisms, with and without decoherence effects in $3+1$ formalism varying $a$ from 0 to 11 and keeping the neutrino energy 11 PeV and baseline of $10^{12}$ km.
The flavour compositions of the neutrinos $\nu_{e}$, $\nu_{\mu}$, $\nu_{\tau}$, and $\nu_{s}$ are clearly illustrated in a normalized form in Fig. 8 and Tab.  5. The post-oscillation observations, which exclude decoherence effects and assume $\nu_{s}=0$ at the source, deliver the compositions: (0.33 : 0.34 : 0.25 : 0.08), (0.43 : 0.27 : 0.24 : 0.05), (0.65 : 0.09 : 0.22 : 0.06), and (0.13 : 0.51 : 0.27 : 0.9) for the four initial compositions outlined in Sec. 6.6, respectively. However, when we incorporate the effects of decoherence, the data showcases a significant shift in these values. The revised compositions are (0.34 : 0.34 : 0.24 : 0.08), (0.40 : 0.30 : 0.22 : 0.08), (0.58 : 0.19 : 0.19 : 0.04), and (0.20 : 0.42 : 0.27 : 0.09), showcasing finite differences in both cases. This evidence decisively indicates that the inclusion of heavy sterile neutrinos has a profound impact on flavour composition measurements on Earth, and advancing decoherence studies will undoubtedly yield critical insights into these scenarios.
@@ 8 Summary and Conclusion
After briefly discussing the existing treatments of neutrino oscillations, we show the mathematical formalism of describing neutrinos as an open quantum system, where we discuss the possibility of neutrino decoherence in the three and $3+1$ flavour scenarios. We explain the modified dispersion relation and discuss the relevant parameters that contain the vital information regarding the various quantum gravity theories. In this regard, it is to be noted that in order to relate the modified dispersion relation with existing theories, we use a variable $a$ which can take values between 0 and $1$, where $a=1$ denotes that the modification to the dispersion relation is Planck energy suppressed, whereas $a=0$ signifies that the modification is entirely dependent on the energy of the system.
@@
We consider the decoherence parameter (and coherence length), oscillation probability, and the flux composition as the main observables. We have shown that for $a=1$ the decoherence parameter and coherence length are independent of neutrino energy, whereas for $a=0$ the decoherence parameter (coherence length) decreases (increases) with energy. Next we show the evolution of the oscillation probability in this decoherence setup in two different ways. In the first case, we show how the oscillation probabilities change with propagation length and we see that for $a=2/3$ and neutrino energy of $100$ PeV, the oscillation probabilities show a significant change after propagating large distances. In order to understand this behaviour further, we also show the oscillation probabilities averaged over propagation lengths for both normal and inverted hierarchy of neutrino masses. We see that the effect of decoherence manifests differently in normal and inverted hierarchy cases. We then show the evolution of the oscillation probabilities for a $100$ PeV neutrino with in a propagation length ranging from $10^{9}$-$10^{15}$ km and constraining $a$ between 0 and $1$. As expected, the decoherence sets in at very low distances for $a=0$, whereas only for $a=1$ the decoherence is negligible for low energies, but for high energies the decoherence becomes dominant, where the survival (transition) probability sees a drop (rise). Finally, we show the flux composition reaching the Earth and the effect of decoherence in that aspect. We see that the effect of decoherence in this case depends on the initial flux composition, the $a$ value, and the mass hierarchy, i.e., for inverted hierarchy, the deviation of the final flux composition due to decoherence is considerably different than the normal hierarchy.
@@
Apart from the three-flavour case (as discussed above), we also show the effect of decoherence on the oscillation probabilities for a $3+1$ scenario. We show that, however, overall the features of the evolution of survival and transition properties look similar to the three-flavour case, the drop (rise) in the survival (transition) probability has an anomalous drop in the $3+1$ flavour case. This drop can be directly linked to the existence of the extra flavour. On the other hand, due to the presence of the extra flavour, the final flux composition is different in the $3+1$ flavour case than that of the three flavour case.